OpenStudy (anonymous):

how much pure acid must be added to 8mL of 20% acid solution in order to obtain a 50% acid solution?

OpenStudy (anonymous):

So set it up the same way I did the last one. If you run into trouble, ask.

OpenStudy (amistre64):

8(.2) + y(1) = (8+y)(.5) solve for "y" :)

OpenStudy (anonymous):

thats what i came up with! thanks

OpenStudy (anonymous):

ok i tried to solve this and got 1.2 and thats not one of the mulitple choice options what am i doing wrong

OpenStudy (anonymous):

20(8) + 100(x) = 50(8+x) 100x + 160 = 400 + 50x 50x = 400 - 160 5x = 24 x = 4.8 ml